# Kudzu: Fast and Simple High-Throughput BFT

### v1

@@ 1. Introduction
    Integrated slow path. If more than $p$ replicas fail, Kudzu maintains the best possible finalization latency of $3\delta$ by running a slow path alongside the fast path.
@@
Technical Intuition. Inspired by DispersedSimplex (song, ), Kudzu introduces the minimal changes to that protocol required to incorporate a fast path. A designated leader in Kudzu distributes the block data by sending erasure coded fragments to all other replicas. In turn, these replicas then broadcast the fragments themselves, together with a first-round vote on the cryptographic hash identifying the block. The voting logic carefully incorporates a rule that makes sure that a block can already be finalized given $n-p$ first-round votes. As a result, replicas can already reassemble the block and finalize it in $2\delta$ latency in the good case where the network is well behaved, the leader is honest, and at most $p$ replicas are corrupt. Another vote-counting rule is used to finalize the block on a slower, $3\delta$-latency path, if more than $p$ replicas are corrupt. To avoid getting stuck altogether, a replica may also vote for a couple of other blocks than the one it initially received from the leader, including a special “timeout” block. The main innovation of our paper is the logic for determining when these “extra” votes are cast — this logic is deceptively simple, but carefully maintains a very delicate balance between liveness and safety. With this innovation, we avoid intricacies such as “progress certificates”, complex view changes, and speculative execution as found in other protocols (such as Banyan (banyan, ), SBFT (gueta2019sbft, ), Kuznetsov et al. (kuznetsov2021revisiting, ), and HotStuff-1 (kang2024hotstuff, )).
@@ 2.1. Fast Path Protocols
A long line of work proposes consensus protocols with a fast path, typically called “fast”, “early-stopping” or “one-step” consensus (brasileiro2001consensus, ; kursawe2002optimistic, ; friedman2005simple, ; song2008bosco, ; martin2006fast, ; guerraoui2007refined, ). FaB Paxos (martin2006fast, ) introduces a parametrized model with $13f+2p+$ 1 replicas, where $p\geq 0$. The parameter $p$ describes the number of replicas that are not needed for the fast path. These protocols can terminate optimally fast in theory ($2\delta$, or 2 network delays) under optimistic conditions. The paper (martin2006fast, ) proves a corresponding lower bound of $13f+2p+$ 1 on the number of replicas needed to achieve this fast behavior. However, as pointed out by Kuznetsov et al. (kuznetsov2021revisiting, ) and Abraham et al. (abraham2021good, ), this lower bound actually only applies to a restricted type of protocol. In fact, (kuznetsov2021revisiting, ) presents a single-shot consensus protocol that uses only $13f+2p^{*}-$ 1 replicas, with $p^{*}\geq 1$, and proves a corresponding lower bound.
2.2. Integrated Slow Path Protocols
Interestingly, in practice, these fast path protocols might increase the finalization latency, as the fast path requires a round of voting between $n-p$ replicas, which could be slower than two rounds of voting between $n-f-p$ replicas that are concentrated in a geographic area. Banyan (banyan, ) performs the fast path in parallel with the $3\delta$ mechanism, which is optimally fast if more than $p$ replicas are faulty or exhibit higher network latency. However, Banyan can exhibit unbounded message complexity when there is a corrupt leader. Kudzu addresses this drawback, shares the same optimistic latency properties, improves throughput through balanced dispersal, benefits from a simpler design, and better properties during leader rotation (see section 2.4).
@@ 2.3. High-Throughput Protocols
Leader-based protocols such as PBFT (PBFT-short, ), HotStuff (HotStuff, ) and Tendermint (Tendermint, ) suffer from a bandwidth bottleneck, as in these protocols the leader is responsible for disseminating transactions to all replicas. As mentioned above, this bottleneck has been well known for quite a while (MirBFT, ; HoneyBadger, ; Narwhal, ). One way to alleviate this bottleneck is to move away from leader-based protocols to a more symmetric, leaderless protocol design where all replicas disseminate transactions. Such an approach was already taken in (MirBFT, ; HoneyBadger, ; DagRider, ; Narwhal, ). These protocols typically have worse latency than leader-based protocols, and moreover, since many replicas may end up broadcasting the same transactions, the supposed improvement in throughput can end up being illusory (note that (MirBFT, ) actually tackles this duplication problem head on). Protocols such as HotStuff (HotStuff, ) and SBFT (gueta2019sbft, ) aim to eliminate the all-to-all broadcasts that occur in the voting logic of many other atomic broadcast protocols, but in so doing increase the finalization latency of the protocol while doing nothing to eliminate the leader bottleneck.
A different approach to address leader bottlenecks makes use of erasure coding (solana, ). By splitting blocks into smaller, erasure-coded shares, the leader can transmit less data, leading to a balanced utilization of resources. This line of work shines by providing high throughput and low latency (song, ; yang2022dispersedledger, ).
@@ 2.4. View Changes
Some atomic broadcast protocols are notorious for having complex view changes, especially in the presence of a fast path (aublin2015next, ). In the case of Zzyzzva (kotla2007zyzzyva, ) and UpRight (clement2009upright, ), safety errors were later pointed out (abraham2017revisiting, ). Improved protocols such as Thelma, Velma and Zelma (abraham2018zelma, ) revisit and correct these.
@@ 3. Model and Preliminaries
We consider a network of $n$ replicas $P_{1},\dots,P_{n}$ called replicas. Up to $f$ replicas can be Byzantine, i.e., deviate from the protocol in arbitrary ways, such as collude to attack the protocol. The remaining replicas follow the protocol and are referred to as honest. We aim to provide better latency if only up to $p$ replicas do not cooperate. In other words, $n-p$ honest replicas including the leader are enough for the fast path to be effective.
@@ 3.1. Network Assumptions
We will not generally assume network synchrony. However, we say the network is $\delta$-synchronous at time $T$ if every message sent from an honest replica $P$ at or before time $T$ to an honest replica $Q$ is received by $Q$ before time $T+\delta$. We also say the the network is $\delta$-synchronous over an interval $[a,b+\delta]$ if it is $\delta$-synchronous at time $T$ for all $T\in[a,b]$.
@@ 3.2. Problem Statement
  • Fast Termination: If the network is in a period of synchrony, $n-p$ replicas behave momentarily honestly, and an honest leader proposes a block $B$ at time $t$, then every honest replica finalizes $B$ at time $t+2\delta$.
@@ 3.3. Cryptographic Assumptions
As we will see, we will need a one $k$-out-of-$n$ threshold signature scheme with $k=n-f-p$, and another with $k=n-p$.
@@
We make use of well-known techniques for asynchronous verifiable information dispersal (AVID) involving erasure codes and Merkle trees(DBLP:conf/wdag/CachinT05a, ).
@@
Under collision resistance for the hash function used for the Merkle trees, any $f+p+1$ certified fragments for given tag $\tau$ will decode to the same payload — moreover, if $\tau$ is the output of the encoding function, these fragments will decode to $M$ (and therefore, if the decoding function outputs $\bot$, we can be sure that $\tau$ was maliciously constructed). This observation is the basis for the protocols in (SodsBC, ; DBLP:conf/podc/LuL0W20, ; DBLP:journals/corr/abs-2110-04371, ). Moreover, with this approach, we do not need to use anything like an “erasure code proof system” (as in (cryptoeprint:2021/1500, )), which would add significant computational complexity (and in particular, the erasure coding would have to be done using parameters compatible with the proof system, which would likely lead to much less efficient encoding and decoding algorithms).
@@ 4.1. Protocol Data Objects
We also call a certified fragment for the tag $\tau$ a certified fragment for $B$.
The block $B_{v}^{\textsf{timeout}}=\textsf{Block}(v,\bot,\bot)$ is a special timeout block.
Votes and certificates.
@@
A notarization vote from $P_{i}$ for block $B$ is an object of the form $\textsf{NotarVote}(B,\sigma_{i},f_{i},\pi_{i})$, where $\sigma_{i}$ is a valid signature share from $P_{i}$ on the object $\textsf{Notar}(B)$, and $(f_{i},\pi_{i})$ is either a certified fragment for $B$ at position $i$, or $(f_{i},\pi_{i})=(\bot,\bot)$ if $B=B_{v}^{\textsf{timeout}}$. A notarization certificate for $B$ is an object of the form $\textsf{NotarCert}(B,\sigma)$, where $\sigma$ is a valid $(n-f-p)$-out-of-$n$ signature certificate on the object $\textsf{Notar}(B)$.
The notarization vote on the timeout block is also called the timeout vote, and the notarization certificate for the timeout block is called the timeout certificate.
A first vote from $P_{i}$ on block $B$ is an object of the form $\textsf{FirstVote}(\sigma_{i}^{\prime},\textsf{NotarVote}(B,\sigma_{i},f_{i},\pi_{i}))$, where $\sigma_{i}^{\prime}$ is a valid signature share from $P_{i}$ on the object $\textsf{First}(B)$, and $\textsf{NotarVote}(B,\sigma_{i},f_{i},\pi_{i})$ is a notarization vote from $P_{i}$ on block $B$. A fast finalization certificate for $B\neq B_{v}^{\textsf{timeout}}$ is an object of the form $\textsf{FirstCert}(B,\sigma)$, where $\sigma$ is a valid $(n-p)$-out-of-$n$ signature certificate on the object $\textsf{First}(B)$.
A finalization vote from $P_{i}$ on block $B$ is an object of the form $\textsf{FinalVote}(B,\sigma_{i})$, where $\sigma_{i}$ is a valid signature share from $P_{i}$ on the object $\textsf{Final}(B)$. A finalization certificate for $B$ is an object of the form $\textsf{FinalCert}(v,\sigma)$, where $\sigma$ is a valid $(n-f-p)$-out-of-$n$ signature certificate on the object $\textsf{Final}(B)$.
4.2. Vote and Certificate Pool.
@@
Whenever a replica receives enough votes, and it does not already have a corresponding certificate, it will generate the certificate, add it to the pool, and broadcast the certificate to all replicas. Similarly, whenever a replica receives a certificate, and it does not already have a corresponding certificate, it will add it to the pool, and broadcast the certificate to all replicas. For one slot it is impossible that the pool would receive or create more than: 1 timeout certificate, 1 fast finalization certificate, 1 finalization certificate, and 5 notarization certificates. The first bound is immediate, since there can be only one timeout certificate per round. The second and third bounds follow from the safety analysis below. The fourth bound follows from the analysis in Section 5.3.
4.3. Complete Block Tree.
@@ 4.4. Finalization
4.5. Generating Block Proposals.
The logic for generating block proposal material $B$, $(f_{1},\pi_{1}),\ldots,(f_{n},\pi_{n})$ in slot $v$ in line 17 of Protocol 1 is as follows:
@@
4.6. Validating Block Proposals.
To check if $\textsf{BlockProp}(B,f_{j},\pi_{j})$ is a valid block proposal from the leader in slot $v$ in line 20 of Protocol 1, replica $P_{j}$ checks that the following conditions holds:
@@
  • $(f_{j},\pi_{j})$ is a certified fragment for $\tau$ at position $j$.
Additionally, we require the following conditions to also be checked:
@@ 4.7. Main Loop
[…]
For example, if the pool contains 1, 2, 3, 4 first votes on blocks $B_{1}$, $B_{2}$, $B_{3}$, $B_{v}^{\textsf{timeout}}$ respectively (and $f+p+1=2$), then $\textsf{allVotes}(v)=10$, $\textsf{maxVotes}(v)=3$, and $\textsf{manyVotes}(v)=\{B_{2},B_{3}\}$.
The protocol also uses a subprotocol ReconstructAndNotarize$(v,B)$, defined in Protocol 2.
@@
  • Lines 810 present the logic for the replica successfully exiting the slot by finding a block $B$ for that slot in its complete block tree. In addition, if the replica did not broadcast a notarization vote for any other block (including the timeout block) for that slot, it will also broadcast a finalization block for $B$.
  • Lines 1213 present the logic for the replica unsuccessfully exiting the slot by obtaining a timeout certificate for that slot.
  • Lines 1518 present the logic for the replica proposing a block for that slot if it is the leader for that slot. It generates the block proposal as in Section 4.5, extending the path in the complete block tree ending at $B_{\text{\rm p}}$. Here, $B_{\text{\rm p}}$ is either the genesis block or the block that it found in its complete block tree the last time it successfully exited a slot (other choices of $B_{\text{\rm p}}$ are possible).
  • Lines 2023 present the logic for the replica broadcasting a first vote for a non-timeout block $B$. It will do so only if $B$ is a valid block proposed by the leader (as in Section 4.6) and has not already first voted. Recall that a first vote for $B$ also includes a notarization vote for $B$.
  • Lines 2528 present the logic for the replica broadcasting a first vote for the non-timeout block for this slot. It will do so only if a sufficient amount of time has passed since it entered the slot and has not already first voted.
[…]
  • Lines 3336 present the logic for the replica taking a “second look” at a block $B$, if it has received sufficiently many first votes for $B$. It will do so only if it has already first voted (and has not already taken a second look at $B$). If it can reconstruct a valid payload for $B$, it will broadcast a notarization vote for $B$ (if it has not already done so); otherwise, it will broadcast a notarization vote for the timeout block.
  • Lines 3841 present the logic for the replica broadcasting a timeout vote under special circumstances. It will do so only if it has already first voted and
@@
2:$T_{\text{\rm start}}\leftarrow\textsf{clock}()$triangleright slot‑local initialisation
3:$\textit{done},\textit{proposed},\textit{firstVoted}\leftarrow\mathsf{false}$
4:$\textit{notarized}\leftarrow\{\}$triangleright blocks already notarized
5:$\textit{secondLook}\leftarrow\{\}$triangleright blocks already reconsidered
6:$\textit{firstVote}\leftarrow\{\}$triangleright map $P_{i}\mapsto B$ for their first vote
@@
8:there exists a block $B$ for slot $v$ in the complete block tree $\Rightarrow$
9:     $B_{\text{\rm p}}\leftarrow B$; $\textit{done}\leftarrow\mathsf{true}$
10:     if $\textit{notarized}\subseteq\{B\}$ then broadcast $\textsf{FinalVote}(B,\sigma_{j})$
11:
12:the pool contains a timeout certificate for $v$Rightarrow
13:     $\textit{done}\leftarrow\mathsf{true}$
14:
15:$\lnot\textit{proposed}\;\land\;\textsf{leader}(v)=P_{j}\Rightarrow$
16:     $\textit{proposed}\leftarrow\mathsf{true}$
17:     generate block proposal material $B$, $(f_{1},\pi_{1}),\dots,(f_{n},\pi_{n})$ extending block $B_{\text{\rm p}}$
18:     for all $i\in[n]$: send $\textsf{BlockProp}(B,f_{i},\pi_{i})$ to $P_{i}$
19:
[…]
21:     $\textit{firstVoted}\leftarrow\mathsf{true}$
22:     broadcast $\textsf{FirstVote}(\sigma_{j}^{\prime},\textsf{NotarVote}(B,\sigma_{j},f_{j},\pi_{j}))$
23:     $\textit{notarized}\leftarrow\textit{notarized}\cup\{B\}$
24:
25:$\lnot\textit{firstVoted}\;\land\;\textsf{clock}()>T_{\text{\rm start}}+\Delta_{\text{\rm timeout}}\Rightarrow$
26:     $\textit{firstVoted}\leftarrow\mathsf{true}$
27:     broadcast $\textsf{FirstVote}(\sigma_{j}^{\prime},\textsf{NotarVote}(B_{v}^{\textsf{timeout}},\sigma_{j},\bot,\bot))$
28:     $\textit{notarized}\leftarrow\textit{notarized}\cup\{B_{v}^{\textsf{timeout}}\}$
29:
[…]
31:     $\textit{firstVote}[P_{i}]\leftarrow B$
32:
33:$\textit{firstVoted}\;\land\;\exists B\in\textsf{manyVotes}(\textit{firstVote})\setminus\textit{secondLook}$
[…]
35:     $\textit{secondLook}\leftarrow\textit{secondLook}\cup\{B\}$
36:     ReconstructAndNotarize$(v,B)$
37:
38:firstVoted(allVotes(firstVote)maxVotes($\textit{firstVoted}\;\land\;(\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\geq f+p+1)$
[…]
40:     broadcast $\textsf{NotarVote}(B_{v}^{\textsf{timeout}},\sigma_{j},\bot,\bot)$
[…]
Protocol 2 ReconstructAndNotarize$(v,B)$
[…]
2:broadcast $\textsf{NotarVote}(B,\sigma_{j},f_{j},\pi_{j})$
[…]
4:broadcast $\textsf{NotarVote}(B_{v}^{\textsf{timeout}},\sigma_{j},\bot,\bot)$
[…]

### v2

@@ 1. Introduction
    Simultaneous slow path. If more than $p$ replicas fail, Kudzu maintains the best possible finalization latency of $3\delta$ by running a slow path alongside the fast path.
@@
Technical Intuition. Inspired by DispersedSimplex (song, ), Kudzu introduces the minimal changes to that protocol required to incorporate a fast path. A designated leader in Kudzu distributes the block data by sending erasure coded fragments to all other replicas. In turn, these replicas then broadcast the fragments themselves, together with a first-round vote on the cryptographic hash identifying the block. The voting logic carefully incorporates a rule that makes sure that a block can already be finalized given npn-p first-round votes. As a result, replicas can already reassemble the block and finalize it in $2\delta$ latency in the good case where the network is well behaved, the leader is honest, and at most $p$ replicas are corrupt. Another vote-counting rule is used to finalize the block on a slower, $3\delta$-latency path, if more than $p$ replicas are corrupt. To avoid getting stuck altogether, a replica may also vote for a couple of other blocks than the one it initially received from the leader, including a special “timeout” block. The main innovation of our paper is the logic for determining when these “extra” votes are cast — this logic is deceptively simple, but carefully maintains a very delicate balance between liveness and safety. With this innovation, we avoid intricacies such as “progress certificates”, complex view changes, and speculative execution as found in other protocols (such as Banyan (banyan, ), SBFT (gueta2019sbft, ), Kuznetsov et al. (kuznetsov2021revisiting, ), and HotStuff-1 (kang2024hotstuff, )).
@@ 2.1. Fast Path Protocols
A long line of work proposes consensus protocols with a fast path, typically called “fast”, “early-stopping” or “one-step” consensus (brasileiro2001consensus, ; kursawe2002optimistic, ; dutta2004complexity, ; song2008bosco, ; martin2006fast, ; guerraoui2007refined, ). FaB Paxos (martin2006fast, ) introduces a parametrized model with $3f+2p+1$ replicas, where $p\geq 0$. The parameter $p$ describes the number of replicas that are not needed for the fast path. These protocols can terminate optimally fast in theory ($2\delta$, or 2 network delays) under optimistic conditions. The papers (dutta2004complexity, ; kuznetsov2021revisiting, ; abraham2021good, ) point out that the lower bound of $3f+2p+1$ actually only applies to a restricted type of protocol. The papers present single-shot consensus protocols that use only $3f+2p^{*}-1$ replicas, with $p^{*}\geq 1$, and prove the corresponding lower bound.
2.2. Simultaneous Slow Path Protocols
Interestingly, in practice, these fast path protocols might increase the finalization latency, as the fast path requires a round of voting between npn-p replicas, which could be slower than two rounds of voting between nfpn-f-p replicas that are concentrated in a geographic area. Banyan (banyan, ) performs the fast path in parallel with the $3\delta$ mechanism, which is optimally fast if more than $p$ replicas are faulty or exhibit higher network latency. However, Banyan can exhibit unbounded message complexity when there is a corrupt leader. Kudzu addresses this drawback, shares the same optimistic latency properties, improves throughput through balanced dispersal, benefits from a simpler design, and better properties during leader rotation (see section 2.4).
@@ 2.3. High-Throughput Protocols
Leader-based protocols such as PBFT (PBFT-short, ), HotStuff (HotStuff, ) and Tendermint (Tendermint, ) suffer from a bandwidth bottleneck, as in these protocols the leader is responsible for disseminating transactions to all replicas. As mentioned above, this bottleneck has been well known for quite a while (MirBFT, ; HoneyBadger, ; Narwhal, ). One way to alleviate this bottleneck is to move away from leader-based protocols to a more symmetric, leaderless protocol design where all replicas disseminate transactions. Such an approach was already taken in (MirBFT, ; HoneyBadger, ; DagRider, ; Narwhal, ). These protocols typically have worse latency than leader-based protocols, and moreover, since many replicas may end up broadcasting the same transactions, the supposed improvement in throughput can end up being illusory (note that (MirBFT, ) actually tackles this duplication problem head on).
Kudzu follows a different approach to address the leader bottleneck, which makes use of erasure coding (solana, ; turbine, ). By splitting blocks into smaller, erasure-coded shares, the leader can transmit less data, leading to a balanced utilization of resources.
This line of work shines by providing high throughput and low latency (song, ; pipes, ). Alpenglow (alpenglow, ) is a recent proof-of-stake protocol being deployed for the Solana blockchain that showcases the appeal of this line of work in practice. Alpenglow and Kudzu are intimately related in their voting logic.
@@ 2.4. View Changes
Some atomic broadcast protocols are notorious for having complex view changes, especially in the presence of a fast path (aublin2015next, ). In the case of Zzyzzva (kotla2007zyzzyva, ) and UpRight (clement2009upright, ), safety errors were later pointed out (abraham2017revisiting, ). Arguably SBFT (gueta2019sbft, ) was the first to correct these mistakes. SBFT (gueta2019sbft, ) avoids the all-to-all broadcast that occurs in Kudzu, but this results in higher finalization latency. Moreover, SBFT does not address the above-mentioned leader bottleneck.
@@ 3. Model and Preliminaries
We consider a network of $n$ replicas $P_{1},\dots,P_{n}$ called replicas. Up to $f$ replicas can be Byzantine, i.e., deviate from the protocol in arbitrary ways, such as collude to attack the protocol. The remaining replicas follow the protocol and are referred to as honest. We aim to provide better latency if only up to $p$ replicas do not cooperate. In other words, npn-p honest replicas including the leader are enough for the fast path to be effective.
@@ 3.1. Network Assumptions
We will not generally assume network synchrony. However, we say the network is $\delta$-synchronous at time $T$ if every message sent from an honest replica $P$ at or before time $T$ to an honest replica $Q$ is received by $Q$ before time $T+\delta$. We also say the the network is $\delta$-synchronous over an interval $a,b+\delta]$ if it is $\delta$-synchronous at time $T$ for all $T\in[a,b]$.
@@ 3.2. Problem Statement
  • Fast Termination: If the network is in a period of synchrony, npn-p replicas behave momentarily honestly, and an honest leader proposes a block $B$ at time $t$, then every honest replica finalizes $B$ at time $t+2\delta$.
@@ 3.3. Cryptographic Assumptions
As we will see, we will need a one $k$-out-of-$n$ threshold signature scheme with k=nfpk=n-f-p, and another with k=npk=n-p.
@@
We make use of well-known techniques for asynchronous verifiable information dispersal (AVID) involving erasure codes and Merkle trees (avid, ).
@@
Under collision resistance for the hash function used for the Merkle trees, any $f+p+1$ certified fragments for given tag $\tau$ will decode to the same payload — moreover, if $\tau$ is the output of the encoding function, these fragments will decode to $M$ (and therefore, if the decoding function outputs $\bot$, we can be sure that $\tau$ was maliciously constructed). This observation is the basis for the protocols in (SodsBC, ; DBLP:conf/podc/LuL0W20, ; yang2022dispersedledger, ). Moreover, with this approach, we do not need to use anything like an “erasure code proof system” (as in (cryptoeprint:2021/1500, )), which would add significant computational complexity (and in particular, the erasure coding would have to be done using parameters compatible with the proof system, which would likely lead to much less efficient encoding and decoding algorithms).
@@ 4.1. Protocol Data Objects
We also call a certified fragment for the tag $\tau$ a certified fragment for $B$. The block $B_{v}^{\textsf{timeout}}=\textsf{Block}(v,\bot,\bot)$ is a special timeout block.
@@
A notarization vote from $P_{i}$ for block $B$ is an object of the form $\textsf{NotarVote}(B,\sigma_{i},f_{i},\pi_{i})$, where $\sigma_{i}$ is a valid signature share from $P_{i}$ on the object $\textsf{Notar}(B)$, and $(f_{i},\pi_{i})$ is either a certified fragment for $B$ at position $i$, or $(f_{i},\pi_{i})=(\bot,\bot)$ if $B=B_{v}^{\textsf{timeout}}$.
A notarization certificate for $B$ is an object of the form $\textsf{NotarCert}(B,\sigma)$, where $\sigma$ is a valid (nfp)(n-f-p)-out-of-$n$ signature certificate on the object $\textsf{Notar}(B)$. The notarization vote on the timeout block is also called the timeout vote, and the notarization certificate for the timeout block is called the timeout certificate.
A first vote from $P_{i}$ on block $B$ is an object of the form $\textsf{FirstVote}(\sigma_{i}^{\prime},\textsf{NotarVote}(B,\sigma_{i},f_{i},\pi_{i}))$, where $\sigma_{i}^{\prime}$ is a valid signature share from $P_{i}$ on the object $\textsf{First}(B)$, and $\textsf{NotarVote}(B,\sigma_{i},f_{i},\pi_{i})$ is a notarization vote from $P_{i}$ on block $B$.
A fast finalization certificate for $B\neq B_{v}^{\textsf{timeout}}$ is an object of the form $\textsf{FirstCert}(B,\sigma)$, where $\sigma$ is a valid (np)(n-p)-out-of-$n$ signature certificate on the object $\textsf{First}(B)$.
A finalization vote from $P_{i}$ on block $B$ is an object of the form $\textsf{FinalVote}(B,\sigma_{i})$, where $\sigma_{i}$ is a valid signature share from $P_{i}$ on the object $\textsf{Final}(B)$.
A finalization certificate for $B$ is an object of the form $\textsf{FinalCert}(B,\sigma)$, where $\sigma$ is a valid (nfp)(n-f-p)-out-of-$n$ signature certificate on the object $\textsf{Final}(B)$.
4.2. Vote and Certificate Pool
@@
Whenever a replica receives enough votes, and it does not already have a corresponding certificate, it will generate the certificate, add it to the pool, and broadcast the certificate to all replicas. Similarly, whenever a replica receives a certificate, and it does not already have a corresponding certificate, it will add it to the pool, and broadcast the certificate to all replicas. For one slot it is impossible that the pool would receive or create more than: 1 timeout certificate, 1 fast finalization certificate, 1 finalization certificate, and 5 notarization certificates. The first bound is immediate, since there can be only one timeout certificate per slot. The second and third bounds follow from the safety analysis below. The fourth bound follows from the analysis in Section 5.3.
4.3. Complete Block Tree
@@ 4.4. Finalization
4.5. Generating Block Proposals
The logic for generating block proposal material $B$, $(f_{1},\pi_{1}),\ldots,(f_{n},\pi_{n})$ in slot $v$ in line 19 of Protocol 1 is as follows:
@@ 4.5. Generating Block Proposals
4.6. Validating Block Proposals
To check if $\textsf{BlockProp}(B,f_{j},\pi_{j})$ is a valid block proposal in slot $v$ in line 22 of Protocol 1, replica $P_{j}$ checks that the following conditions holds:
  • the proposal is signed or otherwise authenticated by the leader,
@@ 4.6. Validating Block Proposals
  • $(f_{j},\pi_{j})$ is a certified fragment for $\tau$ at position $j$,
@@ 4.7. Main Loop
As mentioned in Section 4.2, each replica only considers only one first vote that it receives from any other replica. To make this explicit in the protocol, we use a map $\mathit{firstVote}$ from replicas to blocks to record these votes for the current slot. The protocol also uses simple helper functions on this map.
  • $\textsf{allVotes}(\textit{firstVote})$: the total number of first votes for slot $v$ contained in the pool,
  • $\textsf{maxVotes}(\textit{firstVote})$: the maximal number of first votes on some non-timeout block $B$ for slot $v$ contained in the pool,
  • $\textsf{manyVotes}(\textit{firstVote})$: returns the set of non-timeout blocks in slot $v$ on which the pool contains at least $f+p+1$ first votes.
For example, if the pool contains 1, 2, 3, 4 first votes on blocks $B_{1}$, $B_{2}$, $B_{3}$, $B_{v}^{\textsf{timeout}}$ respectively (and $f+p+1=2$), then $\textsf{allVotes}=10$, $\textsf{maxVotes}=3$, and $\textsf{manyVotes}=\{B_{2},B_{3}\}$.
The protocol also uses a subprotocol ReconstructAndNotarize(v,B)(v,B), defined in Protocol 2.
@@
  • Lines 1012 present the logic for the replica successfully exiting the slot by finding a block $B$ for that slot in its complete block tree. In addition, if the replica did not broadcast a notarization vote for any other block (including the timeout block) for that slot, it will also broadcast a finalization block for $B$.
  • Lines 1415 present the logic for the replica unsuccessfully exiting the slot by obtaining a timeout certificate for that slot.
  • Lines 1720 present the logic for the replica proposing a block for that slot if it is the leader for that slot. It generates the block proposal as in Section 4.5, extending the path in the complete block tree ending at $B_{\text{\rm p}}$. Here, $B_{\text{\rm p}}$ is either the genesis block or the block that it found in its complete block tree the last time it successfully exited a slot (other choices of $B_{\text{\rm p}}$ are possible).
  • Lines 2225 present the logic for the replica broadcasting a first vote for a non-timeout block $B$. It will do so only if $B$ is a valid block proposed by the leader (as in Section 4.6) and has not already first voted. Recall that a first vote for $B$ also includes a notarization vote for $B$.
  • Lines 2730 present the logic for the replica broadcasting a first vote for the timeout block for this slot. It will do so only if a sufficient amount of time has passed since it entered the slot and has not already first voted.
  • Lines 3233 present the logic for updating the map $\mathit{firstVotes}$. No other actions are taken.
  • Lines 3538 present the logic for the replica taking a “second look” at a block $B$, if it has received sufficiently many first votes for $B$. It will do so only if it has already first voted (and has not already taken a second look at $B$). If it can reconstruct a valid payload for $B$, it will broadcast a notarization vote for $B$ (if it has not already done so); otherwise, it will broadcast a notarization vote for the timeout block.
  • Lines 4043 present the logic for the replica broadcasting a timeout vote under special circumstances. It will do so only if it has already first voted and
@@
1: $B_{\text{\rm p}}\leftarrow\textsf{genesis}$ $\triangleright$ parent of the next block
2: for $v=1,2,\dots$ do
3:   $T_{\text{\rm start}}\leftarrow\textsf{clock}()$ $\triangleright$ slot‑local initialisation
4:   $\textit{done},\textit{proposed},\textit{firstVoted}\leftarrow\mathsf{false}$
5:   $\textit{notarized}\leftarrow\{\}$ $\triangleright$ blocks already notarized
6:   $\textit{secondLook}\leftarrow\{\}$ $\triangleright$ blocks already reconsidered
7:   $\textit{firstVote}\leftarrow\{\}$ $\triangleright$ map $P_{i}\mapsto B$ for their first vote 8:
9:   while $\lnot\textit{done}$ wait until either
10:    there exists a block $B$ for slot $v$ in the complete block tree $\Rightarrow$
11:    $B_{\text{\rm p}}\leftarrow B$; $\textit{done}\leftarrow\mathsf{true}$
12:
if $\textit{notarized}\subseteq\{B\}$ then broadcast $\textsf{FinalVote}(B,\sigma_{j})$
13:
14:    the pool contains a timeout certificate for $v$ $\Rightarrow$
15:    $\textit{done}\leftarrow\mathsf{true}$ 16:
17:    $\lnot\textit{proposed}\;\land\;\textsf{leader}(v)=P_{j}\Rightarrow$
18:    $\textit{proposed}\leftarrow\mathsf{true}$
19:    generate block proposal material $B$, $(f_{1},\pi_{1}),\dots,(f_{n},\pi_{n})$ extending block $B_{\text{\rm p}}$
20:    for all $i\in[n]$: send $\textsf{BlockProp}(B,f_{i},\pi_{i})$ to $P_{i}$ 21:
22:    $\lnot\textit{firstVoted}\;\land\;$received valid $\textsf{BlockProp}(B,f_{j},\pi_{j})$ from $\textsf{leader}(v)\Rightarrow$
23:    $\textit{firstVoted}\leftarrow\mathsf{true}$
24:    broadcast $\textsf{FirstVote}(\sigma_{j}^{\prime},\textsf{NotarVote}(B,\sigma_{j},f_{j},\pi_{j}))$
25:    $\textit{notarized}\leftarrow\textit{notarized}\cup\{B\}$ 26:
27:    $\lnot\textit{firstVoted}\;\land\;\textsf{clock}()>T_{\text{\rm start}}+\Delta_{\text{\rm timeout}}\Rightarrow$
28:    $\textit{firstVoted}\leftarrow\mathsf{true}$
29:    broadcast $\textsf{FirstVote}(\sigma_{j}^{\prime},\textsf{NotarVote}(B_{v}^{\textsf{timeout}},\sigma_{j},\bot,\bot))$
30:    $\textit{notarized}\leftarrow\textit{notarized}\cup\{B_{v}^{\textsf{timeout}}\}$ 31:
@@
33:    $\textit{firstVote}[P_{i}]\leftarrow B$ 34:
35:    $\textit{firstVoted}\;\land\;\exists B\in\textsf{manyVotes}(\textit{firstVote})\setminus\textit{secondLook}$
36:    andBB’s parent is in the complete block tree $\Rightarrow$
37:    $\textit{secondLook}\leftarrow\textit{secondLook}\cup\{B\}$
38:    ReconstructAndNotarize(v,B)(v,B) 39:
40:    $\textit{firstVoted}\;\land\;(\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\geq f+p+1)$
@@
42:    broadcast $\textsf{NotarVote}(B_{v}^{\textsf{timeout}},\sigma_{j},\bot,\bot)$
43:    $\textit{notarized}\leftarrow\textit{notarized}\cup\{B_{v}^{\textsf{timeout}}\}$
Protocol 2 ReconstructAndNotarize(v,B)(v,B)
1: reconstruct payload for $B$
2: if reconstruction succeeds $\land$ payload valid then
3:   if $B\notin\textit{notarized}$ then
4:    broadcast $\textsf{NotarVote}(B,\sigma_{j},f_{j},\pi_{j})$
5:    $\textit{notarized}\leftarrow\textit{notarized}\cup\{B\}$
6: else
7:   if $B_{v}^{\textsf{timeout}}\notin\textit{notarized}$ then
8:    broadcast $\textsf{NotarVote}(B_{v}^{\textsf{timeout}},\sigma_{j},\bot,\bot)$
9:    $\textit{notarized}\leftarrow\textit{notarized}\cup\{B_{v}^{\textsf{timeout}}\}$